OpenStudy (anonymous):

Use Laplace transform to solve the following system: x''=x+3y y''=4x-(4e^t) x(0)=2; x'(0)=3 ;y(0)=1 ;y'(0)=2

OpenStudy (turingtest):

Laplace to solve a system? new one on me, can't wait to see the soution

OpenStudy (turingtest):

...but you know how to take the laplace of these expressions, right?

OpenStudy (anonymous):

yeah but i have no idea how to solve this system D=

OpenStudy (turingtest):

me neither :( let me call for reinforcements: @JamesJ @Mr.Math @across @lalaly help with a DE system

OpenStudy (anonymous):

transform both DEs\[s^{2} X(s) - sx(0) - x'(0) = X(s) + 3Y(s)\]\[s^{2}Y(s) - sy(0) - y'(0) = 4X(s)-\frac{4}{s-1}\]plug in the initial conditions, simplify and solve for X(s) and Y(s)

OpenStudy (mr.math):

It's easy if you already know how to do the laplace transform. You'll be left, as exraven explained, with two equations in two variables \(X(s)\) and \(Y(s)\), solve for them and then take the inverse laplace transform to each one separately.